# TopoFR: A Closer Look at Topology Alignment on Face Recognition

### v1

@@ 4.2 Structure Damage Estimation
In practical FR scenarios, low-quality face samples, also known as "hard samples", are commonly included in the training set. These hard samples tend to be encoded in abnormal positions near the decision boundary in the latent space [25, 6, 62, 26], which will disrupt the latent space’s topological structure and further hinder the alignment of structures. To address this issue, we propose a novel hard sample mining strategy called Structure Damage Estimation (SDE). SDE is specifically designed to identify hard samples with serious structure damage within the training set accurately. By prioritizing the learning of these hard samples and guiding them back to the reasonable positions during optimization, SDE aims to mitigate the adverse impact of hard samples on the latent space’s topological structure.
Prediction Uncertainty. Hard samples are typically distributed near the decision boundary, thus have a high prediction uncertainty (i.e., large entropy of the classifier prediction) and are more likely to be misclassified by the classifier. Conversely, easy samples are usually located far from the decision boundary and have relatively low prediction uncertainty. To model the difficulty of each sample, we introduce a binary random variable $u_{i}\in\left\{0,1\right\}$ for each sample $\widetilde{x}_{i}$ to indicate whether the sample is hard or easy by values of 1 and 0, respectively. Then the probability that sample $\widetilde{x}_{i}$ belongs to hard samples (i.e., with large prediction uncertainty) can be defined as $h_{\varphi}(\widetilde{x}_{i})=P_{\varphi}\left(u_{i}=1|\widetilde{x}_{i}\right)$, where $\varphi$ represents the parameter set. According to the cluster assumption [63], we believe that samples with higher prediction entropy are more disruptive to the latent space’s topological structure. Therefore, we propose to model the distribution of the entropy $E(\widetilde{g}_{i})$ for each training sample $\widetilde{x}_{i}$ using the Gaussian-uniform mixture (GUM) model, a statistical distribution that is robust to outliers [64, 65, 66]:
@@
Assume $\widehat{E}(\widetilde{g}_{i})=(-1)^{\epsilon_{i}}E(\widetilde{g}_{i})$, $\epsilon_{i}\sim B(1,0.5)$, where $B$ is a Bernoulli distribution [67], then the variable $\widehat{E}(\widetilde{g}_{i})$ obeys the following statistical distribution:
@@
In this way, the maximum likelihood model of Equation (6) can be formulated as: $\max\limits_{\pi,\Sigma,\Omega}\displaystyle\prod_{i=1}^{n}p\left(\widehat{E}(\widetilde{g}_{i})|\widetilde{x}_{i}\right)$. Then, the parameter set $\varphi=\left\{\pi,\Sigma,\Omega\right\}$ of GUM can be estimated via the Expectation-Maximization (EM) algorithm [68] with the following iterative formulas:
@@
Specially, at each iteration, EM alternates between evaluating the expected log-likelihood (E-step) and updating the parameter set $\varphi$ (M-step). In Equation (7), the E-step aims to evaluate the posterior probability $h_{\varphi}^{(t+1)}$ of an sample $\widetilde{x}_{i}$ to be hard sample using the iterative formula $h_{\varphi}^{(t+1)}(\widetilde{x}_{i})$, where $(t+1)$ denotes the EM iteration index. The M-step updates the parameter set $\varphi$ using the iterative formulas $\pi^{(t+1)}$, $\Sigma^{(t+1)}$ and $\Omega^{(t+1)}$.
Structure Damage Score (SDS). Compared to correctly classified samples, misclassified samples usually have larger difficulty and have greater destructive effects on the latent space’s topological structure. Therefore, misclassified samples need to receive more attention during training. Inspired by the Focal loss [69], we design a probability-aware scoring mechanism $\omega(\widetilde{x}_{i})$ that combines prediction uncertainty $h_{\varphi}(\cdot)$ and prediction accuracy to adaptively compute SDS for each sample $\widetilde{x}_{i}$:
@@ 5 Experiments
i) For training, we employ three distinct datasets, namely MS1MV2 [1] (5.8M facial images, 85K identities), Glint360K [38] (17.1M facial images, 360K identities), and WebFace42M [70] dataset (42.5M facial images, 2M identities). ii) For evaluation, we adopt LFW [71], AgeDB-30 [72], CFP-FP [73], CPLFW [74], CALFW [75], IJB-C [23], IJB-B [76] and the ICCV-2021 Masked Face Recognition Challenge (MFR-Ongoing) [27] as the benchmarks to test the performance of our models.
Notably, the MFR-Ongoing [27] is the most authoritative and comprehensive competition for evaluating FR models’ generalization performance. It contains not only the existing popular test sets, such as IJB-C, but also its own MFR benchmarks, such as Mask, Children, and Multi-Racial test sets. Due to page size limitation, more training settings and experimental results are placed on Appendix.
5.2 Results on Mainstream Benchmarks
@@
R200, Partial FC [40] CVPR22 91.87 - 97.79 98.70 98.54 89.52 97.70 96.93 97.97
R200, UniFace [77] WebFace42M ICCV23 92.43 93.11 98.14 98.98 98.84 90.01 97.92 96.68 97.91
@@
R50, ArcFace [1] CVPR19 99.68 97.11 97.53 88.36 92.52 91.66
R50, MagFace [5] CVPR21 99.74 97.47 97.70 88.95 93.34 91.47
R50, AdaFace [3] CVPR22 99.82 97.86 97.85 - 96.27 94.42
@@
R100, CosFace [2] CVPR18 99.78 98.26 98.17 92.68 95.56 94.01
R100, ArcFace [1] CVPR19 99.77 98.27 98.15 92.69 95.74 94.09
R100, MV-Softmax [78] AAAI20 99.80 98.28 97.95 - 95.20 93.60
R100, URL [61] CVPR20 99.78 98.64 - 95.00 96.60 -
R100, BroadFace [79] ECCV20 99.85 98.63 98.38 94.59 96.38 94.97
MS1MV2 R100, CurricularFace [4] CVPR20 99.80 98.37 98.32 - 96.10 94.80
R100, MagFace+ [5] CVPR21 99.83 98.46 98.17 94.08 95.97 94.51
R100, SCF-ArcFace [25] CVPR21 99.82 98.40 98.30 94.04 96.09 94.74
R100, DAM-CurricularFace [80] ICCV21 - - - - 96.20 95.12
R100, ElasticFace-Cos+ [81] CVPR22 99.80 98.73 98.28 - 96.65 95.43
R100, AdaFace [3] CVPR22 99.82 98.49 98.05 - 96.89 95.67
TransFace-B [9] ICCV23 99.82 98.39 98.27 94.15 96.55 -
@@
R200, ArcFace [1] CVPR19 99.79 98.44 98.19 94.67 96.53 95.18
R200, AdaFace [3] CVPR22 99.83 98.76 98.28 94.88 96.93 95.71
TransFace-L [9] ICCV23 99.83 98.65 98.23 94.55 96.59 -
@@
R50, ArcFace [1] CVPR19 99.78 98.77 98.28 95.29 96.81 95.30
R50, AdaFace [3] CVPR22 99.82 99.07 98.34 95.58 96.90 95.66
@@
R100, ArcFace [1] CVPR19 99.81 99.04 98.31 95.38 96.89 95.69
R100, AdaFace [3] CVPR22 99.82 99.20 98.58 96.24 97.19 95.87
Glint360K TransFace-B [9] ICCV23 99.85 99.17 98.53 96.18 97.45 -
@@
R200, ArcFace [1] CVPR19 99.82 99.14 98.49 95.71 97.20 95.89
R200, AdaFace [3] CVPR22 99.83 99.24 98.61 95.96 97.33 96.12
TransFace-L [9] ICCV23 99.85 99.32 98.62 96.29 97.61 -
@@
Results on LFW, CFP-FP and AgeDB-30. We adopt MS1MV2 and Glint360K to train our models, respectively. The results are reported in Table 2. To showcase the universality of our method, we also provide detailed experimental results of TopoFR model trained by CosFace [2]. As stated in Refs.[3, 5], the performances of existing FR models on these three benchmarks have reached saturation. 1) On MS1MV2 training set, we note that our TopoFR and TopoFR models still obtain accuracy improvement and outperform SOTA competitors (e.g., AdaFace [3] and TransFace [9]). 2) On Glint360K training set, our TopoFR become SOTA models and surpass AdaFace and TransFace.
@@ A.1 Implementation Details
Training Details. For MS1MV2 and Glint360K, our models are trained using Pytorch on 4 NVIDIA Tesla A100 GPUs, and a mini-batch of 128 images is assigned for each GPU. In the case of WebFace42M, we train our models (ResNet-200 backbone) using 64 NVIDIA Tesla V100 GPUs. We crop all images to 112×112, following the same setting as in ArcFace [1]. For the backbones, we adopt ResNet-50, ResNet-100 and ResNet-200 [82] as modified in [1]. We follow [1] to employ ArcFace ($s=64$ and $m=0.5$) as the basic classification loss to train the TopoFR model. For the TopoFR model trained by CosFace [2], we set the scale $s$ to 64 and the cosine margin $m$ of CosFace to 0.4. To optimize the models, we use Stochastic Gradient Descent (SGD) optimizer with momentum of 0.9 for both datasets. The weight decay for MS1MV2 is set to 5e-4 and 1e-4 for Glint360K. The initial learning rate is set to 0.1 for both datesets. In terms of the balance coefficient $\alpha$, we choose $\alpha=0.1$ for experiments on R50 TopoFR, and $\alpha=0.05$ for experiments on R100 TopoFR and R200 TopoFR. During training, we apply RSP mechanism with a certain probability. Specially, for an original input sample $x$, the probability of it undergoing RSP is $\xi$, and the probability of it remaining unchanged is $1-\xi$. For the hyper-parameter $\xi$, we choose $\xi=0.2$. Notably, in our method, we focus on preserving the 0-dimension homology $H_{0}$ in the topological structure alignment loss $\mathcal{L}_{sa}$. Because preliminary experiments demonstrated that using the 1-dimension or higher-dimension homology only increases model’s training time without clear performance gains. Code and pre-trained models are available at: https://github.com/modelscope/facechain/tree/main/face_module/TopoFR.
@@ A.3 More Ablation Experiments and Analysis
3) Effectiveness of ISA: Previous works often use Bottleneck distance and p-Wasserstein distance to measure the distance between persistence diagrams $\mathcal{D}^{\mathcal{X}}$ and $\mathcal{D}^{\mathcal{\widetilde{Z}}}$ [18, 56]. Concretely, the Bottleneck distance is defined as $\mathcal{L}_{\infty}(\mathcal{D}^{\mathcal{X}},\mathcal{D}^{\mathcal{\widetilde{Z}}})=\inf_{\kappa:\mathcal{D}^{\mathcal{X}}\rightarrow\mathcal{D}^{\mathcal{\widetilde{Z}}}}\sup_{\varpi\in\mathcal{D}^{\mathcal{X}}}\left\|\varpi-\kappa(\varpi)\right\|_{\infty}$, with $\kappa$ ranging over all bijections between sets of persistent intervals in diagrams $\mathcal{D}^{\mathcal{X}}$ and $\mathcal{D}^{\mathcal{\widetilde{Z}}}$, and $\left\|\cdot\right\|_{\infty}$ denotes the $\infty$-∞ -norm. Equivalently, the p-Wasserstein distance is defined as $\mathcal{L}_{p}(\mathcal{D}^{\mathcal{X}},\mathcal{D}^{\mathcal{\widetilde{Z}}})=(\inf_{\kappa:\mathcal{D}^{\mathcal{X}}\rightarrow\mathcal{D}^{\mathcal{\widetilde{Z}}}}\sum_{\varpi\in\mathcal{D}^{\mathcal{X}}}\left\|\varpi-\kappa(\varpi)\right\|^{p}_{\infty})^{1/p}$. However, the Bottleneck distance metric and p-Wasserstein distance metric are sensitive to outliers [51, 57, 86] and will significantly increase the training time of FR models. This makes them unsuitable for FR tasks with extremely large-scale datasets.
@@ 18. Answer: [Yes]
  19. Justification: We have provided the required code and pre-trained model for this paper. Please refer to the anonymous GitHub link provided at the end of the abstract.

### v2

@@ 4.2 Structure Damage Estimation
In practical FR scenarios, low-quality face samples, also known as "hard samples", are commonly included in the training set. These hard samples tend to be encoded in abnormal positions near the decision boundary in the latent space li2021spherical ; chang2020data ; dan2023trust ; liu2024user , which will disrupt the latent space’s topological structure and further hinder the alignment of structures. To address this issue, we propose a novel hard sample mining strategy called Structure Damage Estimation (SDE). SDE is specifically designed to identify hard samples with serious structure damage within the training set accurately. By prioritizing the learning of these hard samples and guiding them back to the reasonable positions during optimization, SDE aims to mitigate the adverse impact of hard samples on the latent space’s topological structure.
Prediction Uncertainty. Hard samples are typically distributed near the decision boundary, thus have a high prediction uncertainty (i.e., large entropy of the classifier prediction) and are more likely to be misclassified by the classifier lu2024improving ; liu2023joint ; liu2023contrastive ; liu2022exploiting . Conversely, easy samples are usually located far from the decision boundary and have relatively low prediction uncertainty. To model the difficulty of each sample, we introduce a binary random variable $u_{i}\in\left\{0,1\right\}$ for each sample $\widetilde{x}_{i}$ to indicate whether the sample is hard or easy by values of 1 and 0, respectively. Then the probability that sample $\widetilde{x}_{i}$ belongs to hard samples (i.e., with large prediction uncertainty) can be defined as $h_{\varphi}(\widetilde{x}_{i})=P_{\varphi}\left(u_{i}=1|\widetilde{x}_{i}\right)$, where $\varphi$ represents the parameter set. According to the cluster assumption chapelle2005semi ; liu2022collaborative , we believe that samples with higher prediction entropy are more disruptive to the latent space’s topological structure. Therefore, we propose to model the distribution of the entropy $E(\widetilde{g}_{i})$ for each training sample $\widetilde{x}_{i}$ using the Gaussian-uniform mixture (GUM) model, a statistical distribution that is robust to outliers de2015using ; lathuiliere2018deepgum ; gu2020spherical :
@@
Assume $\widehat{E}(\widetilde{g}_{i})=(-1)^{\epsilon_{i}}E(\widetilde{g}_{i})$, $\epsilon_{i}\sim B(1,0.5)$, where $B$ is a Bernoulli distribution dai2013multivariate , then the variable $\widehat{E}(\widetilde{g}_{i})$ obeys the following statistical distribution:
@@
In this way, the maximum likelihood model of Equation (6) can be formulated as: $\max\limits_{\pi,\Sigma,\Omega}\displaystyle\prod_{i=1}^{n}p\left(\widehat{E}(\widetilde{g}_{i})|\widetilde{x}_{i}\right)$. Then, the parameter set $\varphi=\left\{\pi,\Sigma,\Omega\right\}$ of GUM can be estimated via the Expectation-Maximization (EM) algorithm coretto2016robust with the following iterative formulas:
@@
Specifically, at each iteration, EM alternates between evaluating the expected log-likelihood (E-step) and updating the parameter set $\varphi$ (M-step). In Equation (7), the E-step aims to evaluate the posterior probability $h_{\varphi}^{(t+1)}$ of an sample $\widetilde{x}_{i}$ to be hard sample using the iterative formula $h_{\varphi}^{(t+1)}(\widetilde{x}_{i})$, where $(t+1)$ denotes the EM iteration index. The M-step updates the parameter set $\varphi$ using the iterative formulas $\pi^{(t+1)}$, $\Sigma^{(t+1)}$ and $\Omega^{(t+1)}$.
Structure Damage Score (SDS). Compared to correctly classified samples, misclassified samples usually have larger difficulty and have greater destructive effects on the latent space’s topological structure. Therefore, misclassified samples need to receive more attention during training. Inspired by the Focal loss lin2017focal , we design a probability-aware scoring mechanism $\omega(\widetilde{x}_{i})$ that combines prediction uncertainty $h_{\varphi}(\cdot)$ and prediction accuracy to adaptively compute SDS for each sample $\widetilde{x}_{i}$:
@@ 5 Experiments
i) For training, we employ three distinct datasets, namely MS1MV2 deng2019arcface (5.8M facial images, 85K identities), Glint360K an2021partial (17.1M facial images, 360K identities), and WebFace42M zhu2022webface260m dataset (42.5M facial images, 2M identities). ii) For evaluation, we adopt LFW huang2008labeled , AgeDB-30 moschoglou2017agedb , CFP-FP sengupta2016frontal , CPLFW zheng2018cross , CALFW zheng2017cross , IJB-C maze2018iarpa , IJB-B whitelam2017iarpa and the ICCV-2021 Masked Face Recognition Challenge (MFR-Ongoing) deng2021masked as the benchmarks to test the performance of our models.
Notably, the MFR-Ongoing deng2021masked is the most authoritative and comprehensive competition for evaluating FR models’ generalization performance. It contains not only the existing popular test sets, such as IJB-C, but also its own MFR benchmarks, such as Mask, Children, and Multi-Racial test sets. Due to page size limitation, more training settings and experimental results are placed on Appendix.
@@
R200, Partial FC an2022killing CVPR22 91.87 - 97.79 98.70 98.54 89.52 97.70 96.93 97.97
R200, UniFace zhou2023uniface WebFace42M ICCV23 92.43 93.11 98.14 98.98 98.84 90.01 97.92 96.68 97.91
@@
R50, ArcFace deng2019arcface CVPR19 99.68 97.11 97.53 88.36 92.52 91.66
R50, MagFace meng2021magface CVPR21 99.74 97.47 97.70 88.95 93.34 91.47
R50, AdaFace kim2022adaface CVPR22 99.82 97.86 97.85 - 96.27 94.42
@@
R100, CosFace wang2018cosface CVPR18 99.78 98.26 98.17 92.68 95.56 94.01
R100, ArcFace deng2019arcface CVPR19 99.77 98.27 98.15 92.69 95.74 94.09
R100, MV-Softmax wang2020mis AAAI20 99.80 98.28 97.95 - 95.20 93.60
R100, URL shi2020towards CVPR20 99.78 98.64 - 95.00 96.60 -
R100, BroadFace kim2020broadface ECCV20 99.85 98.63 98.38 94.59 96.38 94.97
MS1MV2 R100, CurricularFace huang2020curricularface CVPR20 99.80 98.37 98.32 - 96.10 94.80
R100, MagFace+ meng2021magface CVPR21 99.83 98.46 98.17 94.08 95.97 94.51
R100, SCF-ArcFace li2021spherical CVPR21 99.82 98.40 98.30 94.04 96.09 94.74
R100, DAM-CurricularFace liu2021dam ICCV21 - - - - 96.20 95.12
R100, ElasticFace-Cos+ Boutros_2022_CVPR CVPR22 99.80 98.73 98.28 - 96.65 95.43
R100, AdaFace kim2022adaface CVPR22 99.82 98.49 98.05 - 96.89 95.67
TransFace-B dan2023transface ICCV23 99.82 98.39 98.27 94.15 96.55 -
@@
R200, ArcFace deng2019arcface CVPR19 99.79 98.44 98.19 94.67 96.53 95.18
R200, AdaFace kim2022adaface CVPR22 99.83 98.76 98.28 94.88 96.93 95.71
TransFace-L dan2023transface ICCV23 99.83 98.65 98.23 94.55 96.59 -
@@
R50, ArcFace deng2019arcface CVPR19 99.78 98.77 98.28 95.29 96.81 95.30
R50, AdaFace kim2022adaface CVPR22 99.82 99.07 98.34 95.58 96.90 95.66
@@
R100, ArcFace deng2019arcface CVPR19 99.81 99.04 98.31 95.38 96.89 95.69
R100, AdaFace kim2022adaface CVPR22 99.82 99.20 98.58 96.24 97.19 95.87
Glint360K TransFace-B dan2023transface ICCV23 99.85 99.17 98.53 96.18 97.45 -
@@
R200, ArcFace deng2019arcface CVPR19 99.82 99.14 98.49 95.71 97.20 95.89
R200, AdaFace kim2022adaface CVPR22 99.83 99.24 98.61 95.96 97.33 96.12
TransFace-L dan2023transface ICCV23 99.85 99.32 98.62 96.29 97.61 -
@@
5.2 Results on Mainstream Benchmarks
Results on LFW, CFP-FP and AgeDB-30. We adopt MS1MV2 and Glint360K to train our models, respectively. The results are reported in Table 2. To showcase the universality of our method, we also provide detailed experimental results of TopoFR model trained by CosFace wang2018cosface . As stated in Refs.kim2022adaface ; meng2021magface , the performances of existing FR models on these three benchmarks have reached saturation. 1) On MS1MV2 training set, we note that our TopoFR and TopoFR models still obtain accuracy improvement and outperform SOTA competitors (e.g., AdaFace kim2022adaface and TransFace dan2023transface ). 2) On Glint360K training set, our TopoFR become SOTA models and surpass AdaFace and TransFace.
@@ A.1 Implementation Details
Training Details. For MS1MV2 and Glint360K, our models are trained using Pytorch on 4 NVIDIA Tesla A100 GPUs, and a mini-batch of 128 images is assigned for each GPU. In the case of WebFace42M, we train our models (ResNet-200 backbone) using 64 NVIDIA Tesla V100 GPUs. We crop all images to 112×112, following the same setting as in ArcFace [1, 34]. For the backbones, we adopt ResNet-50, ResNet-100 and ResNet-200 [92] as modified in [1]. We follow [1] to employ ArcFace ($s=64$ and $m=0.5$) as the basic classification loss to train the TopoFR model. For the TopoFR model trained by CosFace [2], we set the scale $s$ to 64 and the cosine margin $m$ of CosFace to 0.4. To optimize the models, we use Stochastic Gradient Descent (SGD) optimizer with momentum of 0.9 for both datasets. The weight decay for MS1MV2 is set to 5e-4 and 1e-4 for Glint360K. The initial learning rate is set to 0.1 for both datesets. In terms of the balance coefficient $\alpha$, we choose $\alpha=0.1$ for experiments on R50 TopoFR, and $\alpha=0.05$ for experiments on R100 TopoFR and R200 TopoFR. During training, we apply RSP mechanism with a certain probability. Specially, for an original input sample $x$, the probability of it undergoing RSP is $\xi$, and the probability of it remaining unchanged is $1-\xi$. For the hyper-parameter $\xi$, we choose $\xi=0.2$. Notably, in our method, we focus on preserving the 0-dimension homology $H_{0}$ in the topological structure alignment loss $\mathcal{L}_{sa}$. Because preliminary experiments demonstrated that using the 1-dimension or higher-dimension homology only increases model’s training time without clear performance gains. Code and pre-trained models are available at: https://github.com/modelscope/facechain/tree/main/face_module/TopoFR.
@@ A.3 More Ablation Experiments and Analysis
3) Effectiveness of ISA: Previous works often use Bottleneck distance and p-Wasserstein distance to measure the distance between persistence diagrams $\mathcal{D}^{\mathcal{X}}$ and $\mathcal{D}^{\mathcal{\widetilde{Z}}}$ [18, 60]. Concretely, the Bottleneck distance is defined as $\mathcal{L}_{\infty}(\mathcal{D}^{\mathcal{X}},\mathcal{D}^{\mathcal{\widetilde{Z}}})=\inf_{\kappa:\mathcal{D}^{\mathcal{X}}\rightarrow\mathcal{D}^{\mathcal{\widetilde{Z}}}}\sup_{\varpi\in\mathcal{D}^{\mathcal{X}}}\left\|\varpi-\kappa(\varpi)\right\|_{\infty}$, with $\kappa$ ranging over all bijections between sets of persistent intervals in diagrams $\mathcal{D}^{\mathcal{X}}$ and $\mathcal{D}^{\mathcal{\widetilde{Z}}}$, and $\left\|\cdot\right\|_{\infty}$ denotes the $\infty$-norm. Equivalently, the p-Wasserstein distance is defined as $\mathcal{L}_{p}(\mathcal{D}^{\mathcal{X}},\mathcal{D}^{\mathcal{\widetilde{Z}}})=(\inf_{\kappa:\mathcal{D}^{\mathcal{X}}\rightarrow\mathcal{D}^{\mathcal{\widetilde{Z}}}}\sum_{\varpi\in\mathcal{D}^{\mathcal{X}}}\left\|\varpi-\kappa(\varpi)\right\|^{p}_{\infty})^{1/p}$. However, the Bottleneck distance metric and p-Wasserstein distance metric are sensitive to outliers [55, 61, 96] and will significantly increase the training time of FR models. This makes them unsuitable for FR tasks with extremely large-scale datasets.
@@ 18. Answer: [Yes]
  19. Justification: We have provided the required code and pre-trained model for this paper. Please refer to the GitHub link provided at the end of the abstract.